John used an empty beaker that has a mass of 43.126 g. In a separate jar, he had a mixture of salt and sand. He transferred two scoops of salt and sand mixture to the beaker. The overall mass was 55.449 g. As you know salt is soluble in water, he added enough water to the beaker with salt and sand mixture to completely dissolve the salt. He filtered the mixture and dried what was left in the filter. The mass of that residue was 8.200 g. Based on the experiment calculate:

1. The mass of salt: _____

2. The mass of sand: _____

3. The mass of the mixture (salt + sand): _____




































1
Expert's answer
2022-02-02T06:53:03-0500

1) Mass of salt and sand

"=55.449-43.126=12.323g"


Mass of salt"=12.323-8.200"

"=4.123g"

2)

Mass of sand "=" Mass of residue


"=8.200g"


3)

Mass of the mixture (salt and sand)

"=55.449-43.126"

"=12.323g"